OpenStudy (anonymous):

Simplify the following expression: 2 + 4(4x - 3) My answer is 16x-14

OpenStudy (anonymous):

6(4x - 3) do distribution 24x -18 and there is your answer

OpenStudy (anonymous):

How did you get 24x??

OpenStudy (anonymous):

4 + 2 = what?

OpenStudy (anonymous):

You have to distribute 4 first..

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

not that I remember of..

OpenStudy (misty1212):

your answer is almost right

OpenStudy (misty1212):

but \(2-12=-10\) not \(-14\)

OpenStudy (anonymous):

16x - 18 + 2 16x - 16

OpenStudy (misty1212):

\[2+4(x-3)\\2+4x-12\\ 2x-10\]

OpenStudy (anonymous):

yes, what Misty said.

OpenStudy (anonymous):

sorry did my math wrong :/

OpenStudy (anonymous):

misty is correct

OpenStudy (anonymous):

c:

OpenStudy (anonymous):

2+16x-12 -2 -2 ----------- 16x-14 @misty1212

OpenStudy (anonymous):

You don't subtract, because the 2 isn't over an = sign.

OpenStudy (anonymous):

You only change the sign if it's on the opposite side of an = sign. Such as: 4x -2 = 6 +2 +2 4x = 8 Divide by 4 x=2

OpenStudy (anonymous):

so yes, misty is right.

OpenStudy (anonymous):

Hope this helps. ^-^

OpenStudy (anonymous):

She means 4x -10 though, i think.

OpenStudy (anonymous):

I think @misty1212 had the equation wrong she had it at 2+4(x−3) 2+4x−12 2x−10

OpenStudy (anonymous):

Oh yes, you're right. Hold on.

OpenStudy (anonymous):

So 16x -10..

OpenStudy (anonymous):

@xFrenchFryx isn't -12-2 = -14? so 16x-14?

OpenStudy (anonymous):

Where is the -2 though?

OpenStudy (anonymous):

It's just a 2.

OpenStudy (anonymous):

So it's -12 + 2

OpenStudy (anonymous):

why + 2? here is my solution 2+16x-12 -2 -2 ----------- 16x-14 @xFrenchFryx

OpenStudy (anonymous):

I do have to go though, so I hope you get the answer. ;3

OpenStudy (anonymous):

I explained that you don't change the 2 to a negative unless it's on the opposite side of an = sign.

OpenStudy (anonymous):

Have a good day ^-^